OpenStudy (anonymous):

please someone help me A bungee jumper undergoes simple harmonic motion with amplitude 5.1m and frequency 0.135 Hz. Assume the bungee jumper follows the simple harmonic motion equation x=Acos(wt). Determine the bungee jumper's velocity at 0.25 s. Determine the bungee jumper's velocity at 0.50 s. Determine the bungee jumper's velocity at 1.0 s.

OpenStudy (anonymous):

For this question you know the displacement equation as x=Acos(wt) now if you differentiate a displacement function,with respect to time you get velocity v=dx/dt hence differentiate it you get, v=-Awsin(wt) over here ,you can find velocity be putting the time value ALSO, w=2*pi*n (n=frequency) here place value of frequency to get value of w and put it in the main equation of velocity to get your required solution!!!